Let R be the region bounded by the lines y=2x-1,y=2x-2,y=0 and y=4. By making the substitutions u=(2x-y)/2 ,v=y/2 and integrating over a suitable region evaluate the integral

"\intop"04 "\intop"x=y/2+1 x=y/2. (2x-y)/2 dxdy


1
Expert's answer
2021-07-28T07:24:51-0400

"y=2x-1,y=2x-2,y=0, y=4\\
u=\frac{2x-y}{2}, v=\frac{y}{2} \implies x=u+v, y=2v\\
J= \frac{\partial (x,y)}{\partial (u,v)}=\begin{vmatrix}
 \frac{\partial x}{\partial u} & \frac{\partial y}{\partial u} \\
 \frac{\partial x}{\partial v} & \frac{\partial y}{\partial v}
\end{vmatrix} =\begin{vmatrix}
 1 & 0 \\
 1 & 2
\end{vmatrix} =2"

The integral is "I= \int_0^4 \int_{x=\frac{y}{2}}^{x={\frac{y}{2}}+1}(2x-y) \frac{1}{2}dxdy\\"

"I= \int_{u=0}^{u=1} \int_{v=0}^{v={2}}u|J|dudv\\
I= \int_{0}^{1} \int_{0}^{{2}}u(2)dudv\\
I=2[\frac{u^2}{2}]_0^1[v]_0^2 \\
I=1(2)\\
I=2"